# Emergent Dexterity via Diverse Resets and Large-Scale Reinforcement Learning

### 16 Mar 2026 v1

@@ Abstract
Reinforcement learning in massively parallel physics simulations has driven major progress in sim-to-real robot learning. However, current approaches remain brittle and task-specific, relying on extensive per-task engineering to design rewards, curricula, and demonstrations. Even with this engineering, they often fail on long-horizon, contact-rich manipulation tasks and do not meaningfully scale with compute, as performance quickly saturates when training revisits the same narrow regions of state space. We introduce OmniReset, a simple and scalable framework that enables on-policy reinforcement learning to robustly solve a broad class of dexterous manipulation tasks using a single reward function, fixed algorithm hyperparameters, no curricula, and no human demonstrations. Our key insight is that long-horizon exploration can be dramatically simplified by using simulator resets to systematically expose the RL algorithm to the diverse set of robot-object interactions which underlie dexterous manipulation. OmniReset programmatically generates such resets with minimal human input, converting additional compute directly into broader behavioral coverage and continued performance gains. We show that OmniReset gracefully scales to long-horizon dexterous manipulation tasks beyond the capabilities of existing approaches and is able to learn robust policies over significantly wider ranges of initial conditions than baselines. Finally, we distill OmniReset into visuomotor policies which display robust retrying behavior and substantially higher success rates than baselines when transferred to the real world zero-shot. Project webpage: https://omnireset.github.io
Figure 1: OmniReset automatically generates diverse reset states, enabling complex, multi-phase manipulation behaviors to emerge from large-scale reinforcement learning. Using the same generic reset procedure for each task, OmniReset learns robust, task-specific policies. The first row shows the robot pushing and flipping a drawer before wiggling it in. The second row shows the robot picking up a table leg, adjusting its grip using the table, and twisting the leg into the hole. The final two rows show peg insertion in simulation and a distilled RGB policy attempting the task on a real robot, which recovers from a failed insertion to successfully complete it, demonstrating robust emergent retrying behavior.
@@ 1 Introduction
Reinforcement learning (RL) in massively parallelized simulation environments (Mittal et al., 2023; Todorov et al., 2012) has driven recent successes in sim-to-real robotics (Akkaya et al., 2019; Hwangbo et al., 2019). In principle, these algorithms can automatically acquire complex, contact-rich behaviors through repeated interaction with the environment. Yet in practice, obtaining robust and performant policies still requires extensive per-task engineering, limiting the scalability of this paradigm—particularly for long-horizon manipulation tasks. This sharply contrasts with domains such as language modeling, where simply scaling data and compute has yielded dramatic gains with similarly simple RL algorithms Guo et al. (2025). How can we achieve comparable scalability for robotic manipulation?
@@
The central bottleneck in robotic RL is that standard exploration techniques Schulman et al. (2017); Haarnoja et al. (2018) saturate as parallelism and compute are scaled, repeatedly sampling a narrow state–action distribution and becoming trapped in local minima Singla et al. (2024). While numerous advanced exploration methods have been proposed Pathak et al. (2017); Burda et al. (2018), their added algorithmic complexity makes them difficult to scale in practice. As a result, practitioners often rely on human intuition to reduce the exploration burden, injecting structure through task-specific reward design Westenbroek et al. (2022); Ng and others , hand-designed curricula Lee et al. (2020a), or user-provided demonstrations Bauza et al. (2025); Peng et al. (2018); Nair et al. (2018). Although effective in many settings, these approaches are fundamentally limited by the amount of human effort they require. A complementary strategy is to simplify the learning problem through additional system scaffolding—using RL only for contact-rich phases while relying on motion planning or trajectory optimization for the remainder Lee et al. (2020b); Tang et al. (2024). While this reduces exploration demands, it comes at the cost of increased system complexity. These approaches reflect a prevailing assumption in robotics: that dexterous manipulation is too complex to emerge from large-scale RL alone and must instead be scaffolded with additional task-specific structure.
@@
In contrast, we argue for systematically exposing RL to a superset of the interactions it will encounter when manipulating the scene, and then allowing dexterous task-specific behaviors to emerge from large-scale compute and optimization. Although the space of possible behaviors is vast—flipping, screwing, insertion, and other contact-rich motions—successful policies reuse a relatively small set of recurring interaction modes, such as approaching objects, initiating contact, and forming stable grasps. These modes can be densely covered through generic resets that do not encode task-specific solutions. By sufficiently randomizing object poses and sampling these interaction states, we substantially reduce the exploration burden and ensure the agent encounters meaningful success signals. This coverage allows rewards to propagate smoothly through the state space, enabling the agent to identify high-value regions and learn how to stitch together multiple distinct behaviors to reach these goals, without task-specific shaping our guidance.
@@
Concretely, we introduce OmniReset, a scalable framework for robotic manipulation that automatically generates diverse initial-state distributions to densely cover the contact-rich interactions the robot may encounter. This coverage allows PPO Schulman et al. (2017) to fully leverage large-scale compute, improving performance as the number of parallel environments increases. When scaled sufficiently, OmniReset learns to combine multiple interaction modes—such as pushing, flipping, and insertion (Fig. 1)—into coherent, multi-stage strategies without task-specific reward shaping, curricula, or demonstrations. Across diverse contact-rich tasks, this approach solves long-horizon problems that are out of reach for existing methods, producing robust policies that succeed from a wide range of initial states rather than narrow distributions typical of baseline approaches. Finally, we demonstrate that OmniReset can be used to train visuomotor policies via student-teacher distillation which robustly transfer to the real-world and substantially outperform real-world baselines.
@@ 2 Related Work
Exploiting Resets and in Reinforcement Learning: Exploiting simulator resets for RL is a natural idea which has been explored in many contexts. Prior theoretical works (Kakade and Langford, 2002) have suggested more uniform sampling over initial states, but do not provide practical algorithms. The primary focus of many works is to make learning more tractable by generating an explicit curriculum over resets(Tang et al., 2023; Dennis et al., 2020; Bauza et al., 2025), for instance through a “reverse-curriculum” of states going backwards from the goal (Florensa et al., 2017) or using a learned dynamics model to propose viable resets (Edwards et al., 2018; Ivanovic et al., 2019). In contrast, a second category of methods leverage demonstrations (whether human or automatically generated) to generate feasible pathways to the goal (Tao et al., 2024; Resnick et al., 2018; Salimans and Chen, 2018; Bauza et al., 2025; Tang et al., 2024). In contrast to these prior works, we show that neither human demonstrations, nor a curriculum is needed, but rather that the simple approach for generating diverse resets in OmniReset naturally scales to various long-horizon manipulation problems without added algorithmic complexity.
@@
Leveraging Demonstrations: An alternative is to increasingly rely on human demonstrations and imitation learning to overcome difficult long-horizon exploration. Approaches include adding auxiliary BC loss terms to RL objectives (Nair et al., 2018; Hester et al., 2018; Rajeswaran et al., 2018), simply adding demonstrations to the replay buffer (Vecerik et al., 2017), and introducing reward shaping terms which encourage RL agent to follow demonstrations (Tang et al., 2024; Reddy et al., 2019; Koprulu et al., 2024; Peng et al., 2018). Other works have sought to squeeze more information out demonstrations by automatically translating existing demonstrations to new initial conditions and scenes (Mandlekar et al., 2023) or by robustifying BC policies by promoting recovery behavior (Ke et al., 2023; Ankile et al., 2024). While our work complementarily pushes the limits of what behaviors can be learned entirely from scratch, we expect that demonstrations (when available) can also be incorporated into our framework to further accelerate learning.
@@ 3 Generating Diverse Resets for Learning Dexterous Manipulation
Figure 2: Sim-to-Real Pipeline with OmniReset (1) After generating partial assemblies and grasps from the simulator, (2) we collect reset states: reaching, near object, grasped, and near goal. (3) We then train a state-based RL policy initialized from these reset states, which is used to (4) train student-teacher distillation to get a RGB policy. (5) By finetuning this RGB policy on a mix of simulation data and small set of real demonstrations, (6) we deploy the RGB-based policy in the real world.
This section introduces OmniReset, a scalable framework for robotic manipulation that automatically constructs RL problems by generating diverse, manipulation-centric reset distributions. Rather than relying on task-specific curricula, demonstrations, or carefully tuned rewards to guide exploration, OmniReset exposes standard RL algorithms to a broad superset of key interaction states that would be rarely encountered under naïve exploration. When paired with large-scale simulation and compute, this reset design continually exposes the algorithm to diverse interaction states, preventing convergence to narrow, suboptimal behaviors and enabling complex multi-step manipulation skills to emerge from large-scale optimization.
@@ 3.1 Problem Setting
At a code level, this requires that the user identify $s^{tar}$ in the environment definition, provide a means for sampling goal states, and select and operating region for the robot (such as the area over a table top). These pieces of information place minimal burden on the user, yet nonetheless contain rich information about problem structure which OmniReset exploits to design easily solvable RL problems.
Key Problem Parameters: OmniReset automatically generates the initial state distribution $s_{0}\sim\rho$ with the right diversity which enables PPO to solve each of the tasks we consider with a single task agnostic reward function and zero tuning of algorithm hyper parameters.
@@ 3.2 Automatically Generating RL Problems with Diverse Resets
Reinforcement learning fails to solve long-horizon manipulation tasks when training only encounters a narrow slice of potential object configurations or robot-object interactions. To prevent this collapse, OmniReset systematically expands coverage along two axes. First, we approximately cover the space of pathways along which the target object can be transported to the goal by densely resetting $s^{tar}$ on the tabletop, at random point in the air, and at and near the goal $\mathcal{G}$. This exposes the RL algorithm to difficult-to-discover success signals, and allows these signals to smoothly propagate throughout the state-space as the value function is updated. Next, we cover the different ways the robot can interact with $s^{tar}$ by resetting the arm in configurations where it is reaching towards $s^{tar}$, making contact with $s^{tar}$ from a wide variety of points, and with stable grasps on $s^{tar}$. This exposes the RL algorithm to the different ways the robot can interact with $s^{tar}$ to move it towards the goal. Altogether, this broad coverage enables the RL algorithm to discover high-value regions of the state space and the behaviors required to reach those states.
@@
(a) Grasp Sampling. We display the grasp poses sampled for the table leg. The grasp sampling ranges are broad, moderate, and narrow from left-to-right. Our method uses the broad range.
@@
Altogether, these different reset regions provide dense, approximate coverage over the space of pathways to the goal, without prescribing a priori which behaviors are needed to solve the task. Indeed, as we see from our wide range of examples (Fig.1) the RL algorithm is free to select completely different pathways through the state space when solving different tasks, utilizing the reset states which are useful while ignoring those which are not. For example, in the Drawer Insertion task, at convergence the RL algorithm does not obtain stable grasps on the drawer, but instead repeatedly flips the drawer and then pushes it into the cabinet. In contrast, for the Leg Twisting tasks, the robot picks the table leg, pushes it against the table to obtain a more favorable grasp, and then twists the leg into the hole.
Practical Implementation: In practice it can be difficult to efficiently sample feasible resets which respect the contact constraints of the physics simulator. Sampling invalid initial conditions can lead to pathological and non-physical behavior which destabilize learning. Moreover, reset states must be sampled with minimal overhead to maximize GPU-parallelism. Thus, we first sample feasible resets during an offline phase, sampling proposed resets from the four regions defined above and rejecting invalid samples using a combination of collision checking and stepping the simulator for a few steps to allow object to settle. This yields four validated datasets corresponding to the four reset distributions described above: $D^{R}$, $D^{NO}$, $D^{G}$, and $D^{NG}$. During RL training, we sample from $\mathrm{Uniform}(D)$ where $D=D^{C}\cup D^{NO}\cup D^{G}\cup D^{NG}$. We cache these resets on-GPU to ensure efficient sampling during training.
@@ 3.3 Algorithmic Decisions for RL Training
Next we discuss key algorithmic decisions which are required for PPO to leverage the diversity of resets states and scale to the complexity of tasks we consider.
@@
Here, $r_{\text{success}}$ is a sparse binary reward indicating task completion, $r_{\text{dist}}$ encourages minimizing the distance of the target object $s^{tar}$ to the goal, $r_{\text{reach}}$ encourages the gripper to be near $s^{tar}$, $r_{\text{smooth}}$ penalizes large or rapidly changing actions, and $r_{\text{term}}$ penalizes unsafe or physically invalid states that trigger terminations. Importantly, this reward does not encode task-specific strategies as all components and weights are kept fixed across experiments. We find that this generic structure is sufficient for stable training across diverse manipulation tasks, and performance is largely insensitive to the precise weighting of individual terms. See the Appendix for additional details.
Scaling Parallel Environments: When combined with increasing reset diversity, we found that scaling the number of parallel environments used by PPO stabilized and accelerated learning. This enables OmniReset to obviate curricula and task dependent rewards, significantly reducing the amount of tuning which is required to solve a new task.
Asymmetric Actor-Critic: We asymmetric actor-critic approach (Pinto et al., 2017) for our learning architecture. The actor observations include a history of the five previous time-steps for the state of the robot, the poses of all objects in the scene, and the previous actions taken by the policy. The critic takes in these observations as well as additional privileged parameters of the environment. We found that conditioning the actor on larger observation spaces led to less stable training and led us to only provide this information to the critic.
Generalized State-Dependent Exploration Noise: We employ the policy noise parameterization from (Raffin et al., 2022). In short, gSDE has a separate prediction head which determines the gaussian exploration noise at each time-step and is conditioned on the features of the final layer of the policy network. This approach enables the actor to learn different temporally-correlated exploration strategies in different regions of the state-space, crucial for solving heterogeneous multi-stage tasks.

### 02 Apr 2026 v3

@@ Abstract
Reinforcement learning in massively parallel physics simulations has driven major progress in sim-to-real robot learning. However, current approaches remain brittle and task-specific, relying on extensive per-task engineering to design rewards, curricula, and demonstrations. Even with this engineering, typical reinforcement learning methods can often fail on long-horizon, contact-rich manipulation tasks and do not meaningfully scale with compute, as performance quickly saturates when training revisits the same narrow regions of state space. We introduce OmniReset, a simple and scalable framework that enables on-policy reinforcement learning to robustly solve a broad class of dexterous manipulation tasks using fixed algorithm hyperparameters, no curricula, minimal reward engineering, and no human demonstrations. Our key insight is that long-horizon exploration can be dramatically simplified by using simulator resets to systematically expose the RL algorithm to the diverse set of robot-object interactions that underlie dexterous manipulation. OmniReset programmatically generates such resets with minimal human input, converting additional compute directly into broader behavioral coverage and continued performance gains for dynamic policies. We show that OmniReset gracefully scales to long-horizon dexterous manipulation tasks beyond the capabilities of existing approaches and is able to learn robust policies demonstrating a variety of dynamic, contact-rich recovery behavior. Finally, we distill OmniReset into visuomotor policies that can be transferred to the real world zero-shot, displaying robust retrying behavior to accomplish complex, contact-rich tasks with non-trivial success rates. Project webpage: https://omnireset.github.io
Figure 1: OmniReset is a scalable framework for dexterous manipulation which uses diverse simulator resets and large scale reinforcement learning to solve contact-rich, long horizon tasks beyond the capabilities of existing approaches. These policies are then distilled to RGB camera observations and robustly transferred to the real world zero-shot, where they are able to consistently solve challenging tasks over much wider ranges of initial conditions than baselines.
@@ 1 Introduction
Reinforcement learning (RL) in massively parallelized simulation environments (Mittal et al., 2023; Todorov et al., 2012) has driven recent successes in sim-to-real robotics (Akkaya et al., 2019; Hwangbo et al., 2019). While very successful for locomotion and navigation problems, these methods have seen relatively less success in robotic manipulation problems. In principle, these algorithms can also be applied to robotic manipulation, automatically acquiring complex, contact-rich behaviors through repeated interaction with the simulated environment. Yet in practice, obtaining robust and performant policies still requires extensive per-task environment and reward engineering, limiting the scalability of this paradigm—particularly for long-horizon manipulation tasks. This sharply contrasts with domains such as language modeling, where simply scaling data and compute has yielded dramatic gains with similarly simple RL algorithms Guo et al. (2025). How can we achieve comparable scalability for robotic manipulation?
@@
The central bottleneck in robotic RL is that standard exploration techniques Schulman et al. (2017); Haarnoja et al. (2018) saturate as parallelism and compute are scaled, repeatedly sampling a narrow state–action distribution and becoming trapped in local minima Singla et al. (2024). While numerous advanced exploration methods have been proposed Pathak et al. (2017); Burda et al. (2018), their added algorithmic complexity makes them difficult to scale in practice. As a result, practitioners often rely on human intuition to reduce the exploration burden, injecting structure through task-specific reward design Westenbroek et al. (2022); Ng and others ; Handa et al. (2023), hand-designed curricula Lee et al. (2020a), or user-provided demonstrations Bauza et al. (2025); Peng et al. (2018); Nair et al. (2018). Although effective in many settings, these approaches are fundamentally limited by the amount of human effort they require. A complementary strategy is to simplify the learning problem through additional system scaffolding—using RL only for contact-rich phases while relying on motion planning or trajectory optimization for the remainder Lee et al. (2020b); Tang et al. (2024); Zhou et al. (2024); Lee et al. (2020c). While this reduces exploration demands, it comes at the cost of increased system complexity and a lack of smooth, adaptive behavior for complex tasks. These approaches reflect a prevailing assumption in robotics: that dexterous manipulation is too complex to emerge from large-scale RL alone and must instead be scaffolded with additional task-specific structure.
@@
We argue that with the right system design, this scaffolding is unnecessary. In this work, we instead propose systematically exposing RL to a superset of the interactions it will encounter when manipulating the scene, and then allowing dexterous task-specific behaviors to emerge from large-scale compute and optimization. Although the space of possible behaviors for robust task performance is vast—flipping, screwing, insertion, and other contact-rich motions—successful policies reuse a relatively small set of recurring interaction modes, such as approaching objects, initiating contact, and forming stable grasps. These modes can be densely covered through generic resets that do not encode task-specific solutions. By sufficiently randomizing object poses and sampling these interaction states, we substantially reduce the exploration burden and ensure the agent encounters meaningful success signals. This coverage allows sparse rewards to propagate smoothly through the state space, enabling the agent to identify high-value regions and learn how to stitch together multiple distinct behaviors to reach these goals, without task-specific reward shaping or guidance.
@@
Concretely, we introduce OmniReset, a scalable framework for robotic manipulation that automatically generates diverse initial-state distributions to densely cover the contact-rich interactions the robot may encounter. This coverage allows PPO Schulman et al. (2017) to fully leverage large-scale compute, improving performance as the number of parallel environments increases. When scaled sufficiently, OmniReset learns the emergent behavior of combining multiple interaction modes—such as pushing, flipping, and insertion (Fig. 1)—into coherent, multi-stage strategies without task-specific reward shaping, curricula, or demonstrations. Across diverse contact-rich tasks, this approach solves long-horizon problems that are considerably out of reach for existing methods, producing robust policies that succeed from a huge range of initial states rather than narrow distributions typical of many prior approaches. Finally, we demonstrate that OmniReset can be used to train visuomotor policies via student-teacher distillation which can robustly transfer zero-shot to the real-world, substantially outperforming alternatives such as imitation learning.
@@ 2 Related Work
Exploiting Resets in Reinforcement Learning: Exploiting simulator resets for RL is a natural idea which has been explored in many contexts. Prior theoretical works (Kakade and Langford, 2002) have suggested more uniform sampling over initial states, but do not provide practical algorithms. The primary focus of many works is to make learning more tractable by generating an explicit curriculum over resets (Tang et al., 2023; Dennis et al., 2020; Bauza et al., 2025), for instance through a “reverse-curriculum” of states going backwards from the goal (Florensa et al., 2017) or using a learned dynamics model to propose viable resets (Edwards et al., 2018; Ivanovic et al., 2019). In contrast, a second category of methods leverage demonstrations (whether human or automatically generated) to generate feasible pathways to the goal (Tao et al., 2024; Resnick et al., 2018; Salimans and Chen, 2018; Bauza et al., 2025; Tang et al., 2024). In contrast to these prior works, we show that neither human demonstrations, nor a curriculum is needed, but rather that the simple approach for generating diverse resets in OmniReset naturally scales to various long-horizon manipulation problems without added algorithmic complexity. More recently, staggered environment resets (Bharthulwar et al., 2025) have been shown to improve optimization by decorrelating rollouts, but leave the initial state distribution unchanged and thus do not alleviate exploration. In contrast, OmniReset directly addresses exploration by enumerating diverse, task-relevant initial states, enabling coverage of critical intermediate states that are otherwise rarely visited.
@@
Leveraging Demonstrations: An alternative is to increasingly rely on human demonstrations and imitation learning to overcome difficult long-horizon exploration. Approaches include adding auxiliary BC loss terms to RL objectives (Nair et al., 2018; Hester et al., 2018; Rajeswaran et al., 2017), simply adding demonstrations to the replay buffer (Vecerik et al., 2017), and introducing reward shaping terms which encourage RL agent to follow demonstrations (Tang et al., 2024; Reddy et al., 2019; Koprulu et al., 2024; Peng et al., 2018). Other works have sought to squeeze more information out demonstrations by automatically translating existing demonstrations to new initial conditions and scenes (Mandlekar et al., 2023) or by robustifying BC policies by promoting recovery behavior (Ke et al., 2023; Ankile et al., 2024). While our work complementarily pushes the limits of what behaviors can be learned entirely from scratch, we expect that demonstrations (when available) can also be incorporated into our framework to further accelerate learning.
@@ 3 Generating Diverse Resets for Learning Dexterous Manipulation
Figure 2: Sim-to-Real Pipeline with OmniReset (1) After generating partial assemblies and grasps from the simulator, (2) we collect reset states: reaching, near object, grasped, and near goal. (3) We then train a state-based RL policy initialized from these reset states, which is used to (4) train student-teacher distillation to get a RGB policy. (5) We finally deploy the RGB-based policy in the real world zero-shot.
This section introduces OmniReset, a scalable framework for robotic manipulation that automatically constructs RL problems by generating diverse, manipulation-centric reset distributions in a task-agnostic way. Rather than relying on task-specific curricula, demonstrations, or carefully tuned rewards to guide exploration, OmniReset exposes standard RL algorithms to a broad superset of key interaction states that would be rarely encountered under naïve exploration. When paired with large-scale simulation and compute, this reset design continually exposes the algorithm to diverse interaction states, preventing convergence to narrow, suboptimal behaviors and enabling complex multi-step manipulation skills to emerge from large-scale optimization.
@@ 3.1 Problem Setting
At a code level, this requires that the user identify $s^{tar}$ in the environment definition, provide a means for sampling goal states, and select and operating region for the robot (such as the area over a table top). These pieces of information place minimal burden on the user, yet nonetheless contain rich information about problem structure which OmniReset exploits to design easily solvable RL problems. In short, OmniReset automatically constructs a diverse initial state distribution $s_{0}\sim\rho$ that enables RL algorithms to solve all considered tasks using a single task-agnostic reward function, without any task-specific hyperparameter tuning or explicit curricula.
@@ 3.2 Automatically Generating RL Problems with Diverse Resets
Reinforcement learning fails to solve long-horizon manipulation tasks when training only encounters a narrow slice of potential object configurations or robot-object interactions. To prevent this collapse, OmniReset systematically expands coverage along two axes. First, we approximately cover the space of pathways along which the target object can be transported to the goal by densely resetting $s^{tar}$ on the tabletop, at random point in the air, and at and near the goal $\mathcal{G}$. This exposes the RL algorithm to difficult-to-discover success signals, and allows these signals to smoothly propagate throughout the state-space as the value function is updated. Next, we cover the different ways the robot can interact with $s^{tar}$ by resetting the arm in configurations where it is reaching towards $s^{tar}$, making contact with $s^{tar}$ from a wide variety of points, and with stable grasps on $s^{tar}$. This exposes the RL algorithm to the different ways the robot can interact with $s^{tar}$ to move it towards the goal. Altogether, this broad coverage enables the RL algorithm to discover high-value regions of the state space and the behaviors required to reach those states. In the parlance of motion planning, we are exposing the agent to states in “narrow passages”, allowing them to be easily traversed through RL-based optimization.
@@
(a) Grasp Sampling. We display the grasp poses ($S^{G}\subset S$) sampled for the table leg. From left-to-right, the grasp sampling ranges are broad, moderate, and narrow. Our method uses the broad range.
@@
Altogether, these different reset regions provide dense, approximate coverage over the space of pathways to the goal, without prescribing a priori which behaviors are needed to solve the task. It is important to note two important things about this reset distribution- 1) we do not order or connect the states in any graph structure, the paths between them are completely determined by RL, 2) we do not inform any dynamic behavior between these states, these are completely emergent from the RL process. Indeed, as we see from our wide range of examples (Fig.1), the RL algorithm is free to select completely different pathways through the state space when solving different tasks, utilizing the reset states that are useful while ignoring those that are not. For example, in the Drawer Insertion task, at convergence, the RL algorithm does not obtain stable grasps on the drawer, but instead repeatedly flips the drawer and then pushes it into the cabinet. In contrast, for the Leg Twisting tasks, the robot picks up the table leg, pushes it against the table to obtain a more favorable grasp, and then twists the leg into the hole.
Practical Implementation: In practice, it can be difficult to efficiently sample feasible resets that respect the contact constraints of the physics simulator. Sampling invalid initial conditions can lead to pathological and non-physical behavior that destabilizes learning. Moreover, reset states must be sampled with minimal overhead to maximize GPU-parallelism. Thus, we first sample feasible resets during an offline phase, sampling proposed resets from the four regions defined above and rejecting invalid samples using a combination of collision checking and stepping the simulator for a few steps to allow for stabilization. This yields four validated datasets corresponding to the four reset distributions described above: $D^{R}$, $D^{NO}$, $D^{G}$, and $D^{NG}$. During RL training, we sample from $\mathrm{Uniform}(D)$ where $D=D^{C}\cup D^{NO}\cup D^{G}\cup D^{NG}$. We cache these resets on-GPU to ensure efficient sampling during training.
@@ 3.3 Algorithmic Decisions for RL Training
Next, we discuss key algorithmic decisions that are required for an on-policy algorithm (PPO Schulman et al. (2017)) to learn manipulation behavior by leveraging the diversity of reset states and scale to the complexity of tasks we consider in this work.
@@
Here, $r_{\text{success}}$ is a sparse binary reward indicating task completion, $r_{\text{dist}}$ encourages minimizing the distance of the target object $s^{tar}$ to the goal, $r_{\text{reach}}$ encourages the gripper to be near $s^{tar}$, $r_{\text{smooth}}$ penalizes large or rapidly changing actions, and $r_{\text{term}}$ penalizes unsafe or physically invalid states that trigger terminations. Importantly, this reward does not encode task-specific strategies as all components and weights are kept fixed across experiments. We find that this generic structure is sufficient for stable training across diverse manipulation tasks, and performance is largely insensitive to the precise weighting of individual terms. See Appendix A.1 for additional details.
Scaling Parallel Environments: When combined with increasing reset diversity, we found that scaling the number of parallel environments used by PPO stabilized and accelerated learning. This enables OmniReset to obviate curricula and task dependent rewards, significantly reducing the amount of tuning which is required to solve a new task. Intuitively, this large batch size prevents catastrophic forgetting in situations when many of the reset states result in unsuccessful policy behavior. This is necessary to ensure continued value propagation backwards from the target configuration.
Asymmetric Actor-Critic: Since we have access to privileged information in a simulator, we use an asymmetric actor-critic approach (Pinto et al., 2017) for our learning architecture. The actor observations include a history of the five previous time-steps for the state of the robot, the poses of all objects in the scene, and the previous actions taken by the policy. The critic takes in these observations as well as additional privileged parameters of the environment. We found that conditioning the actor on larger observation spaces led to less stable training and led us to only provide this information to the critic.
Generalized State-Dependent Exploration Noise: We employ the policy noise parameterization from gSDE (Raffin et al., 2022). gSDE has a separate prediction head which determines the gaussian exploration noise at each time-step and is conditioned on the features of the final layer of the policy network. This approach enables the actor to learn different temporally-correlated exploration strategies in different regions of the state-space, crucial for solving heterogeneous multi-stage tasks.